how many ounces of a 10% alcohol solution must be mixed with 18 ounces of a 15% alcohol solution to make a 13% alcohol solution?
Answer by ewatrrr(24785)


10% alcohol solution must be mixed with 18 ounces of a 15% alcohol solution to make a 13% alcohol solution?
*Note: Let x be the amount of the total 10% solution.
.
.10x + .15*18oz = .13(18oz+x)
.15*18oz + .13*18oz = .13x - .10x
.
.02*18oz = .03x
12oz = x
.
check our answer
1.2 + 2.7 = 3.9